Name: Instructor Date: Section 6. The regression line equation for a set of data is given by y-23x+. -10. The value of for this same data set is 10.1. Use α .05 a. If the lincar correlation coefficient for this data is -521, what is the best y -value for x 57 b. If the linear correlation coefficient for this data is r 972, what is the best predicted y-value for x=5? size n 15 with a 01 7. What is the critical value for the linear correlation coefficient r for a sample of 8. The linear correlation coefficient for a set of paired variables (x, y) is r 897. What proportion of the variation in y can be explained by the linear relationship between x and y? 9. identify two different conditions under which the regression line should not be used to make predictions.
Answer #1

6) a) Critical value for n = 10 and a-0.05 is 0.576

Since r<0.576, there is no Significant correlation

So, best predicted value is 10.1

b)  Critical value for n = 10 and a-0.05 is 0.576

Since r>0.576, there is Significant correlation

So, best predicted value is y = 2.3*5+5 = 16.5

7) Critical value for n = 15 and α 0.01 is 0.606

8) r^ 2 = 0.897*0.897 = 0.805

0.805 of variation is explained